Given F(x)= 2x squared-8x+9, find all values of x for which f(x)=9
How do I solve this??

Answer by scott8148(6628)

9 = 2x^2 - 8x + 9

0 = x^2 - 4x

0 = x(x - 4)

x = 0

x - 4 = 0 ___ x = 4